# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ 3.3.4 PIM/NDP
Speculative Decoding. SpecPIM [168] aims to accelerate speculative inference on the PIM-enabled system by extensively exploring the heterogeneity brought by both the algorithm and the architecture. It constructs the architecture design space to satisfy each model’s disparate resource demands and dedicates the dataflow design space to fully utilize the system’s hardware resources. Based on the co-design space, it also proposes a design space exploration framework to provide the optimal design under different target scenarios. Compared with speculative inference on GPUs and existing PIM-based LLM accelerators, SpecPIM achieves 1.52$\times$/2.02$\times$ geomean speedup and 6.67$\times$/2.68$\times$ geomean higher energy efficiency.
3.3.5 Comparison
@@
In Figure 8, for fast decoding, power consumption ranges from 17.499W to 700W, with inference speeds between 24.7 tokens/s and 174.018 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 9.944 tokens/J. SpecPIM [168] (PIM/NDP) achieves the lowest power consumption with 17.499W, the highest throughput with batch size 1, and the highest energy efficiency.
@@
3.4 Operator Optimization
3.4.1 Overview
@@
Figure 9: Operator optimizations.
@@
Table 6: Operator optimization on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.4.2 GPU
Fusion. To address the quadratic memory requirements in the attention computation, FlashAttention [169, 170] fuses the attention operation into a single operator by tiling input matrices (Q, K, V) and the attention matrix into blocks. Based on FlashAttention, FlashDecoding [171] proposes additionally the parallel computation along the feature dimension, improving performance for small batch size during the decoding phase. FlashDecoding achieves 95.07 tokens/s and 54.19 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. Subsequently, FlashDecoding++ [172] optimizes the synchronization overhead in softmax computation by pre-determining a unified maximum based on statistical analysis in advance, enabling the parallel execution of subsequent operations and improving efficiency in typical LLMs like Llama2 and ChatGLM [33]. FlashDecoding++ achieves 115.57 tokens/s and 61.66 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. The linear operator is widely used in deep learning. In the common framework for deep learning (e.g., Pytorch), the backend of the linear operator is usually the General Matrix-Matrix Multiplication (GEMM) operation supported by NVIDIA. The naive implementation by HuggingFace achieves 44.60 tokens/s and 36.29 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. In the LLM framework (e.g., DeepSpeed [173], vLLM [174], and OpenPPL [175]), the GEMM implementation provided by cuBLAS [176] APIs is generally optimized. For the Llama2-7B model on a single NVIDIA A100 GPU, they achieve 89.24 tokens/s, 88.45 tokens/s and 93.63 tokens/s. For the Llama2-7B model on a single NVIDIA RTX 3090 GPU, they achieve 51.28 tokens/s, 53.31 tokens/s and 55.99 tokens/s. To address the inefficiency of GEMM during decoding due to the reduced dimensions, TensorRT-LLM [177] introduces a dedicated General Matrix-Vector Multiplication (GEMV) implementation to support the decoding phase of LLM when batch size equals 1. TensorRT-LLM achieves 98.19 tokens/s for Llama2-7B on a single NVIDIA A100 GPU. For the smaller batch size during the decoding phase, FlashDecoding++ [172] introduces FlatGEMM to address the inefficiencies in cuBLAS [176] and CUTLASS [178] libraries and employs fine-grained tiling and double buffering techniques to improve parallelism and reduce the latency of memory access. Moreover, it adopts a heuristic selection mechanism to dynamically select the most efficient operator based on the input. The performance of FlashDecoding++ is shown above. Operator fusion [169, 179, 173, 172] is a common and effective optimization to reduce the runtime memory access, eliminate kernel launching overhead and enhance parallelism. ByteTransformer [179] and DeepSpeed [173] fuse the main operator (e.g., GEMM) and the lightweight operators(e.g., residual adding, layernorm and activation functions) into a single kernel to reduce the kernel launching overhead. FlashAttention [169] mentioned above fuse the attention operator into one single kernel, reducing significantly the overhead of memory access and the memory requirements. FlashDecoding++ [172] also achieves the integration of seven fused kernels in a transformer block.
3.4.3 ASIC
@@
Fusion. In 2020, Groq company introduces the Tensor Streaming Processor (TSP) architecture [219], a functionally-sliced microarchitecture with memory units interleaved with vector and matrix compute units in order to take advantage of dataflow locality. The first TSP implementation yields a computational density of more than 1 TOPS/$mm^{2}$ for its 25$\times$29 mm 14nm chip at 900 MHz. In 2022, Groq company introduces a novel software-defined communication approach for large-scale interconnection networks of TSP elements [220]. This scalable communication fabric is based on a software-defined dragonfly topology, ultimately yielding a parallel machine learning system with elasticity to support a variety of workloads. Each TSP contributes 220 MB to the global memory capacity, with the maximum capacity limited only by the network’s scale. The large-scale parallel system achieves up to 10,440 TSPs and more than 2 TB of global memory accessible in less than 3ms of end-to-end system latency. Based on two previous works, Groq Language Processing Unit (LPU) [181] is fabricated in 14nm with 185W power consumption. According to [221], LPU achieves 814 tokens/s for Gemma-7B model. Another commercial company, HyperAccel, also proposes a LLM inference chip with dataflow architecture. Latency Processing Unit (LPU) [180] introduces streamlined hardware that maximizes the effective memory bandwidth usage during end-to-end inference regardless of the model size to achieve up to 90% bandwidth utilization for high-speed text generation. It consists of expandable synchronization link (ESL) that hides bulk of the data synchronization latency in a multi-device system to achieve nearperfect scalability. Its on-chip power is 0.284W sythesised by Samsung 4nm and the total system power is 86W with 96GB HBM3. For OPT-1.3B/6.7B/13B/30B, LPU achieves 769.23 tokens/s, 217.39 tokens/s, 112.40 tokens/s and 49.26 tokens/s, respectively. The Wafer-Scale Engine (WSE-3) [188], which powers the Cerebras CS-3 system, is the largest chip ever built. The WSE-3 is 57$\times$ larger than the largest GPU, has 52$\times$ more compute cores, and 880$\times$ more high performance on-chip memory. The only wafer scale processor ever produced, it contains 4 trillion transistors, 900,000 AI-optimized cores, and 44 gigabytes of high performance on-wafer memory. Each wafer consists of 84 dies with 40GB on-chip memory and 97W power consumption. For Llama3.1-8B, it can achieve about 1800 tokens/s, which is 20$\times$ faster than hyperscale clouds.
@@
Coarse-grained Processing. Tensor Contraction Processor (TCP) [184], is composed of coarse-grained PEs, which are designed to be flexible enough to be configured as a large-scale single unit or a set of small independent compute units. TCP chip is designed and fabricated in 5nm technology with 256MB SRAM and 1.5 TB/s 48GB HBM3 under 150W. For Llama2-7B model, TCP achieves about 125 tokens/s with batch size 1 and 1176 tokens/s with batch size 8. Habana Gaudi [185] and Gaudi2 [186] consists of two main compute engines, Matrix Multiplication Engine (MME) and Tensor Processor Core (TPC) cluster. The TPC unit is a SIMD processor tailor-made for general deep learning operations. The biggest difference between GPU and Gaudi architecture is the size of MME. Gaudi architecture can handle 256$\times$256 matrix multiplication, which requires 512 input elements per cycle while GPU architecture with 16$\times$16 Tensor Core requires 8K input elements. Therefore, Gaudi architecture can save 16$\times$ less read bandwidth requirements. Besides compute engine, Gaudi2 includes 96 GB of HBM2E memories delivering 2.45 TB/sec bandwidth, in addition to a 48 MB of local SRAM. Gaudi is fabricated in 16nm and Gaudi2 is fabricated in 7nm technique node with 600W power consumption. For Llama2-7B and Llama2-13B models, Gaudi2 achieves from 81.97 to 111.11 tokens/s and from 49.02 to 64.52 tokens/s with batch size 1, respectively.
3.4.4 PIM/NDP
@@
Nonlinear Function Approximation. AttentionLego [190] proposes a PIM-based matrix-vector multiplication and look-up table-based Softmax design. PIM-GPT [191], which achieves end-to-end acceleration of GPT inference with high performance and high energy efficiency. At the hardware level, PIM-GPT is a hybrid system that includes DRAM-based PIM chips to accelerate VMM near data and an application-specific integrated circuit (ASIC) to support other functions that are too expensive for PIM including necessary nonlinear functions, data communication, and instructions for the PIM chips. At the software level, the mapping scheme is optimized to efficiently support the GPT dataflow. Overall, PIM-GPT achieves 89$\times$ speedup and 221$\times$ energy efficiency over NVIDIA T4 GPU on 8 GPT models with up to 1.4 billion parameters. SAL-PIM [192] also adopts LUT-based linear interpolation to perform complex nonlinear functions in PIM.
Figure 10: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with operator optimization.
Storage Optimization. By observing that a key impediment to truly harness PIM acceleration is deducing optimal data-placement to place the matrix in memory banks, PIMnast [189] proposes matrix tiling/ordering algorithms to tackle these factors and identify orchestration knobs that impact PIM acceleration. For OPT-6.7B model, compared to the SoC baseline (AMD Ryzen PRO 7040 Series processors comprising eight CPU cores, 12 compute units of GPU cores, 16 AIE tiles, and eight channels of LPDDR5 memory for a peak memory bandwidth of 120 GB/s), PIMnast achieves 4.5$\times$ speedup for per-token latency.
3.4.5 Comparison
In Figure 10, for operator optimization, power consumption ranges from 10.44W to 600W, with inference speeds between 4.13 tokens/s and 1800 tokens/s. The energy efficiency ranges from 0.092 tokens/J to 18.556 tokens/J. MARCA [183] (ASIC) achieves the lowest power consumption with 10.44W. Cerebras WSE-3 [188] (ASIC) achieves the highest throughput with batch size 1 (post-silicon results) and the highest energy efficiency.
@@ 3.4.5 Comparison
3.5 Heterogeneous Cooperation
3.5.1 Overview
@@
Figure 11: Heterogeneous cooperation.
@@
Memory Enhancement. focuses on two key aspects: storage capacity enhancement and bandwidth enhancement. As we move into the LLM eras, computational tasks are becoming increasingly complex, leading to greater demands on memory capacity and bandwidth. From the perspective of storage capacity, memory enhancement is achieved by integrating more on-chip memory by using 3D storage stacks on the same chip area. This allows more data and models to be totally stored in a limited on-chip space, reducing the need to access external storage and improving overall efficiency. Another memory enhancement is to accelerate memory-bound computations by increasing data transfer speeds. For heterogeneous cooperation, bandwidth often restricts computational performance. To address this issue, much methods are proposed to support high-speed on-chip [224, 225, 226] and chip-to-chip interconnect, such as Compute Express Link (CXL) [227] and NVLink [63]. These techniques enable low-latency, high-bandwidth data transmission between different chips, significantly speeding up data exchange processes.
Table 7: Heterogeneous cooperation on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.5.2 CPU
Computing Enhancement. (1) CPU-GPU. Only using the CPU may result in slower performance, so many methods employ a combination of CPU and GPU to enhance LLM inference speed. For personal computers, PowerInfer [195] proposes that the hot-activated neurons should be preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators. For various models (OPT-30B/66B, Falcon-40B, and Llama-70B) on a Intel i9-13900K processor equipped with an NVIDIA RTX 4090, it achieves 8.32 tokens/s on average. On a Intel i7-12700K processor equipped with an NVIDIA RTX 2080Ti, it achieves 5.77 tokens/s on average. For smartphones, PowerInfer-2 [196] further leverages the highly heterogeneous XPUs present in smartphone SoCs, such as asymmetric big.LITTLE CPU cores, GPU, and NPU. On OnePlus 12 equipped with SnapDragon 8 Gen 3, PowerInfer-2 achieves 11.7 tokens/s and 10.5 tokens/s on Llama-7B and Mistral-7B, respectively. On OnePlus Ace 2 equipped with SnapDragon 8+ Gen 1, PowerInfer-2 achieves 6.5 tokens/s and 6.3 tokens/s on Llama-7B and Mistral-7B, respectively. For servers, Kim et al. [194] propose an adaptive model to determine the LLM layers to be run on CPU and GPU, respectively, based on the memory capacity requirement and arithmetic intensity. They then propose CPU-GPU cooperative computing that exploits the AMX extensions of the latest Intel CPU. The evaluation demonstrates that for OPT-30B model, CPU-GPU cooperative computing achieves 336 tokens/s with batch size 1 and input tokens 2016 in prefill stage. In decode stage, it achieves about 25 tokens/s with batch size 90 and input tokens 2016, and about 300 tokens/s with batch size 1400 and input tokens 64.
3.5.3 PIM/NDP

### v3

@@ 3.3.4. PIM/NDP
Speculative Decoding. SpecPIM (li2024specpim, ) aims to accelerate speculative inference on the PIM-enabled system by extensively exploring the heterogeneity brought by both the algorithm and the architecture. It constructs the architecture design space to satisfy each model’s disparate resource demands and dedicates the dataflow design space to fully utilize the system’s hardware resources. Based on the co-design space, it also proposes a design space exploration framework to provide the optimal design under different target scenarios. Compared with speculative inference on GPUs and existing PIM-based LLM accelerators, SpecPIM achieves 1.52$\times$/2.02$\times$ geomean speedup and 6.67$\times$/2.68$\times$ geomean higher energy efficiency.
3.3.5. Comparison
@@
In Figure 8, for fast decoding, power consumption ranges from 17.499W to 700W, with inference speeds between 24.7 tokens/s and 174.018 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 9.944 tokens/J. SpecPIM (li2024specpim, ) (PIM/NDP) achieves the lowest power consumption with 17.499W, the highest throughput with batch size 1, and the highest energy efficiency.
@@
3.4. Operator Optimization
3.4.1. Overview
@@
Figure 9. Operator optimizations.
@@
Table 6. Operator optimization on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.4.2. GPU
Fusion. To address the quadratic memory requirements in the attention computation, FlashAttention (flashattention, ; flashattention2, ) fuses the attention operation into a single operator by tiling input matrices (Q, K, V) and the attention matrix into blocks. Based on FlashAttention, FlashDecoding (flashdecoding, ) proposes additionally the parallel computation along the feature dimension, improving performance for small batch size during the decoding phase. FlashDecoding achieves 95.07 tokens/s and 54.19 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. Subsequently, FlashDecoding++ (hong2024flashdecoding, ) optimizes the synchronization overhead in softmax computation by pre-determining a unified maximum based on statistical analysis in advance, enabling the parallel execution of subsequent operations and improving efficiency in typical LLMs like Llama2 and ChatGLM (chatglm, ). FlashDecoding++ achieves 115.57 tokens/s and 61.66 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. The linear operator is widely used in deep learning. In the common framework for deep learning (e.g., Pytorch), the backend of the linear operator is usually the General Matrix-Matrix Multiplication (GEMM) operation supported by NVIDIA. The naive implementation by HuggingFace achieves 44.60 tokens/s and 36.29 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. In the LLM framework (e.g., DeepSpeed (deepspeed, ), vLLM (kwon2023vllm, ), and OpenPPL (OpenPPL, )), the GEMM implementation provided by cuBLAS (cuBLAS, ) APIs is generally optimized. For the Llama2-7B model on a single NVIDIA A100 GPU, they achieve 89.24 tokens/s, 88.45 tokens/s and 93.63 tokens/s. For the Llama2-7B model on a single NVIDIA RTX 3090 GPU, they achieve 51.28 tokens/s, 53.31 tokens/s and 55.99 tokens/s. To address the inefficiency of GEMM during decoding due to the reduced dimensions, TensorRT-LLM (tensorrt-llm, ) introduces a dedicated General Matrix-Vector Multiplication (GEMV) implementation to support the decoding phase of LLM when batch size equals 1. TensorRT-LLM achieves 98.19 tokens/s for Llama2-7B on a single NVIDIA A100 GPU. For the smaller batch size during the decoding phase, FlashDecoding++ (hong2024flashdecoding, ) introduces FlatGEMM to address the inefficiencies in cuBLAS (cuBLAS, ) and CUTLASS (CUTLASS, ) libraries and employs fine-grained tiling and double buffering techniques to improve parallelism and reduce the latency of memory access. Moreover, it adopts a heuristic selection mechanism to dynamically select the most efficient operator based on the input. The performance of FlashDecoding++ is shown above. Operator fusion (flashattention, ; bytetransformer, ; deepspeed, ; hong2024flashdecoding, ) is a common and effective optimization to reduce the runtime memory access, eliminate kernel launching overhead and enhance parallelism. ByteTransformer (bytetransformer, ) and DeepSpeed (deepspeed, ) fuse the main operator (e.g., GEMM) and the lightweight operators(e.g., residual adding, layernorm and activation functions) into a single kernel to reduce the kernel launching overhead. FlashAttention (flashattention, ) mentioned above fuse the attention operator into one single kernel, reducing significantly the overhead of memory access and the memory requirements. FlashDecoding++ (hong2024flashdecoding, ) also achieves the integration of seven fused kernels in a transformer block.
3.4.3. ASIC
@@
Fusion. In 2020, Groq company introduces the Tensor Streaming Processor (TSP) architecture (abts2020think, ), a functionally-sliced microarchitecture with memory units interleaved with vector and matrix compute units in order to take advantage of dataflow locality. The first TSP implementation yields a computational density of more than 1 TOPS/$mm^{2}$ for its 25$\times$29 mm 14nm chip at 900 MHz. In 2022, Groq company introduces a novel software-defined communication approach for large-scale interconnection networks of TSP elements (abts2022software, ). This scalable communication fabric is based on a software-defined dragonfly topology, ultimately yielding a parallel machine learning system with elasticity to support a variety of workloads. Each TSP contributes 220 MB to the global memory capacity, with the maximum capacity limited only by the network’s scale. The large-scale parallel system achieves up to 10,440 TSPs and more than 2 TB of global memory accessible in less than 3ms of end-to-end system latency. Based on two previous works, Groq Language Processing Unit (LPU) (groq_lpu, ) is fabricated in 14nm with 185W power consumption. According to (groq_lpu_result, ), LPU achieves 814 tokens/s for Gemma-7B model. Another commercial company, HyperAccel, also proposes a LLM inference chip with dataflow architecture. Latency Processing Unit (LPU) (moon2024lpu, ) introduces streamlined hardware that maximizes the effective memory bandwidth usage during end-to-end inference regardless of the model size to achieve up to 90% bandwidth utilization for high-speed text generation. It consists of expandable synchronization link (ESL) that hides bulk of the data synchronization latency in a multi-device system to achieve nearperfect scalability. Its on-chip power is 0.284W sythesised by Samsung 4nm and the total system power is 86W with 96GB HBM3. For OPT-1.3B/6.7B/13B/30B, LPU achieves 769.23 tokens/s, 217.39 tokens/s, 112.40 tokens/s and 49.26 tokens/s, respectively. The Wafer-Scale Engine (WSE-3) (lie2024wafer, ), which powers the Cerebras CS-3 system, is the largest chip ever built. The WSE-3 is 57$\times$ larger than the largest GPU, has 52$\times$ more compute cores, and 880$\times$ more high performance on-chip memory. The only wafer scale processor ever produced, it contains 4 trillion transistors, 900,000 AI-optimized cores, and 44 gigabytes of high performance on-wafer memory. Each wafer consists of 84 dies with 40GB on-chip memory and 97W power consumption. For Llama3.1-8B, it can achieve about 1800 tokens/s, which is 20$\times$ faster than hyperscale clouds.
@@
Coarse-grained Processing. Tensor Contraction Processor (TCP) (kim2024tcp, ), is composed of coarse-grained PEs, which are designed to be flexible enough to be configured as a large-scale single unit or a set of small independent compute units. TCP chip is designed and fabricated in 5nm technology with 256MB SRAM and 1.5 TB/s 48GB HBM3 under 150W. For Llama2-7B model, TCP achieves about 125 tokens/s with batch size 1 and 1176 tokens/s with batch size 8. Habana Gaudi (habana_gaudi, ) and Gaudi2 (habana_gaudi2, ) consists of two main compute engines, Matrix Multiplication Engine (MME) and Tensor Processor Core (TPC) cluster. The TPC unit is a SIMD processor tailor-made for general deep learning operations. The biggest difference between GPU and Gaudi architecture is the size of MME. Gaudi architecture can handle 256$\times$256 matrix multiplication, which requires 512 input elements per cycle while GPU architecture with 16$\times$16 Tensor Core requires 8K input elements. Therefore, Gaudi architecture can save 16$\times$ less read bandwidth requirements. Besides compute engine, Gaudi2 includes 96 GB of HBM2E memories delivering 2.45 TB/sec bandwidth, in addition to a 48 MB of local SRAM. Gaudi is fabricated in 16nm and Gaudi2 is fabricated in 7nm technique node with 600W power consumption. For Llama2-7B and Llama2-13B models, Gaudi2 achieves from 81.97 to 111.11 tokens/s and from 49.02 to 64.52 tokens/s with batch size 1, respectively.
3.4.4. PIM/NDP
@@
Nonlinear Function Approximation. AttentionLego (cong2024attentionlego, ) proposes a PIM-based matrix-vector multiplication and look-up table-based Softmax design. PIM-GPT (wu2024pim, ), which achieves end-to-end acceleration of GPT inference with high performance and high energy efficiency. At the hardware level, PIM-GPT is a hybrid system that includes DRAM-based PIM chips to accelerate VMM near data and an application-specific integrated circuit (ASIC) to support other functions that are too expensive for PIM including necessary nonlinear functions, data communication, and instructions for the PIM chips. At the software level, the mapping scheme is optimized to efficiently support the GPT dataflow. Overall, PIM-GPT achieves 89$\times$ speedup and 221$\times$ energy efficiency over NVIDIA T4 GPU on 8 GPT models with up to 1.4 billion parameters. SAL-PIM (han2024sal, ) also adopts LUT-based linear interpolation to perform complex nonlinear functions in PIM.
Figure 10. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with operator optimization.
Storage Optimization. By observing that a key impediment to truly harness PIM acceleration is deducing optimal data-placement to place the matrix in memory banks, PIMnast (ibrahim2024balanced, ) proposes matrix tiling/ordering algorithms to tackle these factors and identify orchestration knobs that impact PIM acceleration. For OPT-6.7B model, compared to the SoC baseline (AMD Ryzen PRO 7040 Series processors comprising eight CPU cores, 12 compute units of GPU cores, 16 AIE tiles, and eight channels of LPDDR5 memory for a peak memory bandwidth of 120 GB/s), PIMnast achieves 4.5$\times$ speedup for per-token latency.
3.4.5. Comparison
In Figure 10, for operator optimization, power consumption ranges from 10.44W to 600W, with inference speeds between 4.13 tokens/s and 1800 tokens/s. The energy efficiency ranges from 0.092 tokens/J to 18.556 tokens/J. MARCA (li2024marca, ) (ASIC) achieves the lowest power consumption with 10.44W. Cerebras WSE-3 (lie2024wafer, ) (ASIC) achieves the highest throughput with batch size 1 (post-silicon results) and the highest energy efficiency.
@@ 3.4.5. Comparison
3.5. Heterogeneous Cooperation
3.5.1. Overview
@@
Figure 11. Heterogeneous cooperation.
@@
Memory Enhancement. focuses on two key aspects: storage capacity enhancement and bandwidth enhancement. As we move into the LLM eras, computational tasks are becoming increasingly complex, leading to greater demands on memory capacity and bandwidth. From the perspective of storage capacity, memory enhancement is achieved by integrating more on-chip memory by using 3D storage stacks on the same chip area. This allows more data and models to be totally stored in a limited on-chip space, reducing the need to access external storage and improving overall efficiency. Another memory enhancement is to accelerate memory-bound computations by increasing data transfer speeds. For heterogeneous cooperation, bandwidth often restricts computational performance. To address this issue, much methods are proposed to support high-speed on-chip (li20204gbps, ; li2023method, ; li2021gain, ) and chip-to-chip interconnect, such as Compute Express Link (CXL) (sharma2022compute, ) and NVLink (foley2017ultra, ). These techniques enable low-latency, high-bandwidth data transmission between different chips, significantly speeding up data exchange processes.
Table 7. Heterogeneous cooperation on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.5.2. CPU
Computing Enhancement. (1) CPU-GPU. Only using the CPU may result in slower performance, so many methods employ a combination of CPU and GPU to enhance LLM inference speed. For personal computers, PowerInfer (song2023powerinfer, ) proposes that the hot-activated neurons should be preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators. For various models (OPT-30B/66B, Falcon-40B, and Llama-70B) on a Intel i9-13900K processor equipped with an NVIDIA RTX 4090, it achieves 8.32 tokens/s on average. On a Intel i7-12700K processor equipped with an NVIDIA RTX 2080Ti, it achieves 5.77 tokens/s on average. For smartphones, PowerInfer-2 (xue2024powerinfer, ) further leverages the highly heterogeneous XPUs present in smartphone SoCs, such as asymmetric big.LITTLE CPU cores, GPU, and NPU. On OnePlus 12 equipped with SnapDragon 8 Gen 3, PowerInfer-2 achieves 11.7 tokens/s and 10.5 tokens/s on Llama-7B and Mistral-7B, respectively. On OnePlus Ace 2 equipped with SnapDragon 8+ Gen 1, PowerInfer-2 achieves 6.5 tokens/s and 6.3 tokens/s on Llama-7B and Mistral-7B, respectively. For servers, Kim et al. (kim2024exploiting, ) propose an adaptive model to determine the LLM layers to be run on CPU and GPU, respectively, based on the memory capacity requirement and arithmetic intensity. They then propose CPU-GPU cooperative computing that exploits the AMX extensions of the latest Intel CPU. The evaluation demonstrates that for OPT-30B model, CPU-GPU cooperative computing achieves 336 tokens/s with batch size 1 and input tokens 2016 in prefill stage. In decode stage, it achieves about 25 tokens/s with batch size 90 and input tokens 2016, and about 300 tokens/s with batch size 1400 and input tokens 64.
3.5.3. PIM/NDP